OpenStudy (anonymous):

0 Real Number Rational Number Irrational Number Integer Whole Number Natural Number

jimthompson5910 (jim_thompson5910):

It's a real number

jimthompson5910 (jim_thompson5910):

It's a rational number, since 0 = 0/1, 0 = 0/2, etc...

jimthompson5910 (jim_thompson5910):

It's an integer

jimthompson5910 (jim_thompson5910):

and it's a whole number

OpenStudy (anonymous):

Thanks :)

jimthompson5910 (jim_thompson5910):

yw